OpenStudy (anonymous):

diff eq. solve equation: (siny)^2dt+tsin2ydy=0

OpenStudy (amistre64):

looks separable to me

OpenStudy (anonymous):

first determine if they're exact. they are. dp/dy=2sinycosy dq/dt=2sinycosy

OpenStudy (amistre64):

might even be exact

OpenStudy (amistre64):

cool lol

OpenStudy (amistre64):

the t sin 2y might be more doable as an integration

OpenStudy (amistre64):

sin(2y) = 2 sin(y) cos(y)

OpenStudy (amistre64):

but then the dt part is just a sdoable lol forgot we could keep the NOTs constant

OpenStudy (anonymous):

my final answer is not the same as the book's so idk if i'm suppose to keep going. i have F=t(siny)^2. book says lnt+2lnsiny=C. when did it become function of e??

OpenStudy (amistre64):

\[sin^2(y)dt+t\ sin(2y)dy=0\] \[\int sin^2(y)dt=t\ sin^2(y)+g(y)\] \[D_y[t\ sin^2(y)+g(y)]=2t\ sin(y)cos(y)+g'(y)\]

OpenStudy (amistre64):

since sin(2y) = 2 sin(y) cos(y) it looks like g'(y) = 0

OpenStudy (anonymous):

yessss

OpenStudy (anonymous):

so F=t(siny)^2y

OpenStudy (amistre64):

g(y) = some arbitrary constant still

OpenStudy (amistre64):

F = t sin^2(y) + C if i read it right

OpenStudy (anonymous):

o ok

OpenStudy (anonymous):

idk why i have in my notes F=function=C...didn't make sense but now i know i prob copied wrong. thanks

OpenStudy (amistre64):

youre welcome, and good luck :)

OpenStudy (anonymous):

ok it still doesn't make sense to me. i did mean to write in my notes function=C and that's the way the book has it too. so back to the question how does it go to C=lnt+2lnsiny? no initial value given

OpenStudy (amistre64):

well, lets see how we do when use the separable method instead of the exact method: \[sin^2(y)\ dt+t\ sin(2y)\ dy=0\] \[t\ sin(2y)\ dy=-sin^2(y)\ dt\] \[\frac{sin(2y)}{-sin^2(y)}\ dy=t^{-1}\ dt\] \[\frac{2sin(y)cos(y)}{-sin(y)sin(y)}\ dy=t^{-1}\ dt\] \[\frac{2cos(y)}{-sin(y)}\ dy=t^{-1}\ dt\] \[\int \left(-2\frac{cos(y)}{sin(y)}\ dy=t^{-1}\ dt\right)\] \[-2ln(sin(y))=ln(t)+C\]

OpenStudy (amistre64):

im sure if I divied that up the other way that Id get +s instead of -s

OpenStudy (anonymous):

thank you amistre, i wish i can figure this out using the exact method tho. good thing about our test coming up is that the instructor will tell us which method to use for each problem...but that can be a bad thing too since we also HAVE to use the specified method

OpenStudy (amistre64):

the separable i believe has an issue when y=0 since we get a 0 in the denominator along the way. the exact does not have that issue: \[sin^2(y)\ dt+t\ sin(2y)\ dy=0\] \[\int sin^2(y)\ dt=t\ sin^2(y)+g(y)\] \[D_y[t\ sin^2(y)+g(y)]=t\ 2sin(y)cos(y)+g'(y)\] \[t\ 2sin(y)cos(y)=t\ sin(2y):\ t\ sin(2y)+g'(y)\] therefore: \[t\ sin(2y)+g'(y)=t\ sin(2y)+ 0 \] \[\int (g'(y)=0)\ \implies \ g(y)=C \] \[y=t\ sin(2y)+C\] and from the separable we get: \[-2ln(sin(y))=ln(t)+C\implies\ \frac{1}{sin^2(y)}=Ce^t\] \[sin^2(y)=Ce^{-t} \implies \ y=sin^{-1}(Ce^{-t/2});\ y\ne 0\]

OpenStudy (amistre64):

as long as you can show how you got to your answer; and define the interval upon which it is defined, you should be good

OpenStudy (anonymous):

thank you thank you